a hemisphere (μ=3) of radius placed on a horizontal surface with its base touching the surface .A  vertical beam of cross sectional radius (√3/2)r is incident symmetrically on its curved  surface.find radius of the spot of light formed on the horizontal surface??????

Grade:12

1 Answers

Askiitians Expert Bharath-IITD
14 years ago

Dear Abhishiek,

We know that For hemissherical surface

f = R/(n-1) and refractive  n= 3 here

and Magnification is given by M= -S2/S1= f/ f-S1 and S2 is on the other side i.e., on the ground we would get +ve otherwise it would have been -ve

By substituting S1= (√3/2)R

We get S2= S1/(√3-1) Where S1= (√3/2)R

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best  !!!

 


Regards,

Askiitians Experts

Adapa Bharath